a wallet contains 20 bills consisting of only ten and five dollar bills. the amount of money in the wallet is $165. how many of each bill is in the wallet?
can you please add the legend?

Answer by macston(5194)

.
F=number of $5 bills; T=number of $10 bills
.
T+F=20
T=20-F Use this to substitute for T below.
.
$5F+$10T=$165 . This converts number of bills to dollars.
$5F+$10(20-F)=$165 . This substitutes for T from above.
$5F+$200-$10F=$165 . Put dollars on right Fs on left and simplify.
-$5F=-$35 . Divide each side by (-$5)
F=7
ANSWER: There are 7 five dollar bills in the wallet.
.
T=20-F=20-7=13
ANSWER 2: There are 13 ten dollar bills in the wallet.
.
CHECK:
$5F+$10T=$165
$5(7)+$10(13)=$165
$35+$130=$165
$165=$165
.